### v1

# Fast-MCS: A Scalable Open-Source Tool to Find Minimal Cut Sets

Abstract
A network is represented as a graph consisting of nodes and edges. A cut set for a source-destination pair in a network is a set of elements that, when failed, cause the source-destination pair to lose connectivity. A MCS (MCS) is a cut set that cannot be further reduced while maintaining its status as a cut set. MCS are crucial in identifying the critical elements in the network that have the most significant impact on failure. This work introduces Fast-MCS, an open-source, scalable tool for evaluating MCS in large, complex networks. Additionally, we compare the computation time of Fast-MCS with the state-of-the-art.
I Introduction
Modern networks, such as communication networks, electrical power grids, transportation, water supply, gas pipelines, etc., can be represented as graphs, consisting of nodes and edges. These networks invariably require high levels of connectivity consistently. A network operator needs to know the most critical elements of the network so they can be better protected in the event of failures.
The connectivity between any source and destination relies on how they are connected in the topology. When a set of elements in the network fails, disconnecting the source-destination pair, that set is called a cut set. A MCS is also a cut set that cannot be further reduced [1]. For example, consider the Germany_17 topology [2] shown in Fig. 1. Let us assume some traffic between the source Ulm and the destination Hamburg. For this source-destination pair, Ulm-Hamburg, (Frankfurt, Leipzig, Hannover) is a cut set because the failure of these nodes causes Ulm and Hamburg to be disconnected. However, the set can be further reduced to (Frankfurt, Leipzig) and still maintain its status as a cut set. Hence, (Frankfurt, Leipzig) is a MCS (MCS). Note that a single source-destination pair can have several MCS of varying sizes. For example, (Dortmund, Hannover, Leipzig) is another MCS of Ulm-Hamburg with size three. Moreover, a node can be part of several MCS. In this example, Leipzig is a part of both the MCS discussed.
Figure 1: Germany_17 topology [2]
MCS are a popular tool used in modern networks for various purposes across various fields, including aerospace, nuclear power, transportation, electrical infrastructure, communication networks, etc. For example, MCS are used in Fault Tree Analysis [3, 4, 5, 6, 7], RBD (RBD) and availability evaluation [8, 9, 10], sovereignty metric evaluation [11, 12], resilience evaluation [13, 14], etc. Note that the network type does not change the concepts discussed in this work. Our Fast-MCS solution is applicable to any network or system that can be represented as a graph consisting of components as nodes and their interconnections as edges. However, this manuscript focuses on communication networks as the primary example.
Several nodes and edges exist in large communication networks. For a source and destination in the network, finding the MCS, i.e., finding the smallest combinations of nodes whose failure can cause loss of connection between that source and destination, gets computationally expensive as the size of the topology increases. This problem is noted to be NP-hard [15]. The aforementioned tools and evaluations depend on finding the MCS first. Therefore, it is essential to find a fast algorithm to evaluate MCS based on the topology information.
Hence, we present the following contributions in this work.
  1. 1.
    Sec. II investigates methods to evaluate MCS.
  2. 2.
    Sec. III discusses our novel method to evaluate find MCS, and its implementation as an open-source tool called Fast-MCS.
  3. 3.
    Sec. IV shows the performance of PyRBD3, and compares it against the state-of-the-art.
II Background and Related Work
The most intuitive method to find MCS is a combinatorial searching algorithm, i.e., checking if each combination of nodes in the network could be a MCS. [16, 8, 17] propose extensions to the classical combinatorial searching algorithm to speed up the process. The authors in [18] extend this approach using Direct Partial Boolean Derivatives (DPBDs). They study the impact of changing a component’s state on the overall system availability, with positive and negative effects on path and cut sets, respectively. However, the aforementioned approaches do not scale well for large topologies because the number of possible node combinations scales as $\sum\limits_{i=1}^{|V|-1}\binom{|V|}{i}$.
Other previous works also consider evaluating MCS. Most of these works discuss the evaluation of MCS based on MPS. A path set between a source and a destination is the complement of a cut set; i.e., a path set is a set of nodes whose availability connects the source and destination. A MPS is also a path set, but it cannot be further reduced. For example, let us assume some traffic between the source Ulm and the destination Hamburg, in the Germany_17 topology in Fig. 1. For this source-destination pair, {Ulm, Muenchen, Nuernberg, Leipzig, Berlin, Hannover, Hamburg} is a path set because the availability of these nodes ensures a successful path between the source and the destination. However, this path set is not minimal because it can be further reduced to {Ulm, Muenchen, Nuernberg, Leipzig, Berlin, Hamburg}. This new set cannot be reduced further, and hence is a MPS. Similar to MCS, a single source-destination pair can have several MPS. For example, {Ulm, Stuttgart, Nuernberg, Leipzig, Berlin, Hamburg} is another MPS. Additionally, a single node can be a part of multiple MPS.
MPS and MCS are the prime implicants of monotone Boolean functions [19] for the success and the failure of a system, respectively. They are also referred to as the minimal satisfying and maximal non-satisfying vectors of a Boolean function [20]. Since a MPS is the logical opposite of a MCS, finding the logical complement of the MPS yields the MCS.
In [21], the authors introduced the INMIN algorithm for inverting path sets and cut sets. However, the approach relies on exhaustive power-set enumeration and does not scale to large systems. Furthermore, the INMIN algorithm only evaluates the cut sets. It does not evaluate MCS. This would require further computation. The authors in [22, 23] introduced a recursive inversion technique based on De Morgan’s laws and Boolean absorption to convert MPS to MCS. Other works, such as [24, 25], use similar approaches to find MCS.
In [26], the authors present their approach based on De Morgan’s rules and Boole-Shannon expansion to evaluate MCS from MPS. Their primary advantage is the ability to split the expression into smaller partitions for faster calculation and recursive absorption. Therefore, the computation required to obtain the minimum-sum-of-products expression of the MCS. Their approach is efficient for smaller networks. However, it does not scale for larger networks. This method is further explained in Sec. III-C.
To overcome the issues identified in previous works, we introduce the following improvements.
  1. 1.
    Our method relies on logically complementing the MPS by using De Morgan’s laws and selects a single common element per step to enable scalability across large topologies.
  2. 2.
    Our method uses a faster, more memory-efficient set-theoretic approach to complement MPS, rather than the long Boolean expressions used in [26].
  3. 3.
    We implement our method as an open-source tool called Fast-MCS [27].
  4. 4.
    We compare the performance of Fast-MCS against the combinatorial method [16, 8, 17] and the Boole-Shannon method [26].
The implementation of the combinatorial method [16, 8, 17] and the Boole-Shannon expansion [26] are explained in Sections III-A and III-C, respectively. Note that this manuscript considers only nodes to facilitate easier explanation. The concepts of MCS and MPS apply to links as well as edges. Furthermore, our tool Fast-MCS can include links in the MCS evaluation, if required.
Figure 2: Combinatorial procedure to evaluate MCS (MCS).
III Methodology
This section discusses the implementation of the combinatorial method [16, 8, 17], the Boole-Shannon expansion [26] technique, and our Fast-MCS. For the Boole-Shannon expansion [26] and Fast-MCS, we use the same algorithm to find MPS.
III-A Combinatorial method to find MCS
The combinatorial method [16, 8, 17] to find the MCS for a source-destination pair is explained in Fig. 2. First, all the simple paths from the source to the destination are noted down as rows. Second, all nodes except the source and destination are treated as potential MCS and listed as columns (in blue). The table is populated as follows.
  1. 1.
    If node $v_{i}$ is used in path $P_{i}$, then the corresponding value is one.
  2. 2.
    If node $v_{i}$ is not used in path $P_{i}$, then the corresponding value is zero.
After populating all the blue columns, if any column has all ones, then the corresponding node is present in all the simple paths from the source to the destination. Therefore, if that node fails, connectivity between the source and the destination will be lost. Hence, this node is a MCS.
Next, all possible two-node combinations are taken as columns (in red). The supersets of previously identified MCS are removed because these supersets can only be cutsets and not MCS. Then, the columns can be populated by using the logical ‘OR’ operation between the respective columns of the nodes in the combination. For example, the column for $v_{i},v_{j}$ is obtained by the logical ‘OR’ operation between the values in the columns for $v_{i}$ and $v_{j}$. After populating the red columns, the test for MCS is performed again by checking the existence of columns with all ones.
This procedure of populating columns for different node combinations and checking for MCS is repeated until no more columns are possible. A collection of all the all-ones columns provides all the MCS for that source-destination pair. For example, in Fig. 2, $(v_{1},v_{2})$ and $(v_{2},v_{3},v_{4})$ are MCS.
III-B Finding MPS (MPS)
We find the MPS based on an extension of the DFS (DFS) algorithm seen in Alg. 1 [9]. First, the branching starts at the source node. The newly explored node is added to the list of visited nodes in that branch. At every step of the branching procedure, two conditions are checked:
  1. 1.
    if the newly explored node is the destination, or
  2. 2.
    if the newly explored node has already been visited in that branch.
If at least one of the aforementioned conditions is true, the branch is terminated. If both conditions are not met, the branching procedure is continued.
Consider Ulm-Hamburg as the example source-destination pair. In the path set {Ulm, Muenchen, Nuernberg, Leipzig, Berlin, Hannover, Hamburg}, Hamburg is reached from Berlin via Hannover. However, Hamburg can be reached directly from Berlin. Therefore, the proposed path set is not minimal. In other words, when a node that can be reached from one of the previously visited nodes through a shorter path is visited through a longer path, that path is no longer minimal.
Hence, the DFS algorithm must be modified to terminate a branch if a new node is being visited that can already be visited from some other node in the visited list directly. This modification to the DFS algorithm, shown in line 10 of Alg. 1, speeds up the search for MPS. The branching procedure continues until all branches have been explored and terminated.
Algorithm 1 Find MPS (MPS) based on the modified DFS (DFS) algorithm.
0: Graph $G=(V,E)$, source srcsrc, destination dstdst.
[…]
1:Call DFS($s$, $\emptyset$)
2:function modified_DFS(current_ncurrent\_n, path_ppath\_p)
3:  Add current_ncurrent\_n to path_ppath\_p.
4:  If current_n=dstcurrent\_n=dst, then
5:   Store path_ppath\_p in $\mathcal{P}$.
6:   Remove current_ncurrent\_n from path_ppath\_p.
7:   return
8:  end if
9:  For each neighbor neighbor_nneighbor\_n of current_ncurrent\_n in $G$, do
10:   If $neighbor\_n\notin path\_p$ and neighbor_nneighbor\_n cannot be reached directly from another node in path_ppath\_p, then
11:    call modified_DFS(neighbor_nneighbor\_n, path_ppath\_p)
12:   end if
13:  end for
14:  Remove current_ncurrent\_n from path_ppath\_p.
15:  return
Figure 3: Example topology, source-destination pair, and its corresponding MPS (MPS).
Figure 4: Fast-MCS’s algorithm illustrated with the example source-destination pair, STS-T, in Fig. 3.
III-C Finding MCS (MCS) based on Boole-Shannon expansion
In this method, the system’s success, i.e., the source-destination pair’s successful connectivity, is represented as its MPS. Failure is represented by Boolean complementation [26]. A pivot element that appears most frequently is chosen to apply the Boole-Shannon expansion. Shannon expansion is applied as follows.
$S=xS_{x=1}+\bar{x}S_{x=0},$ (1)
where $x$ is the pivot element. The $S_{x=1}$ and $S_{x=0}$ expressions are recursively split using the Boole-Shannon expansion, until there are no unique pivot elements. The resulting final expression can be simplified and checked for absorption to obtain the MCS. For example, in the first step, considering the example shown in Fig. 3, for the source-destination pair S-T, the success function is,
$S=AB+ADEF+CDEF+CDB.$ (2)
If $D$ is the pivot element, the next step splits $S$ as,
$S_{D=1}=AB+AEF+CEF+CB,$ (3)
$S_{D=0}=AB.$ (4)
Then the expressions are complemented and considered for further expansion. The final expression is reduced by Boolean absorption to obtain the MCS.
III-D Finding MCS (MCS) using Fast-MCS
Algorithm 2 Fast-MCS’s MCS (MCS) computation using a Decision Tree
0: Set of minimal paths $\mathcal{P}$ from $S$ to $T$
0: Set of minimal cut sets $\mathcal{M}$
1: Remove $S$ and $T$ from each path in $\mathcal{P}$
[…]
3: Begin Decision Tree with $\mathcal{C}$ as the root node $\mathcal{R}$. Call $\mathcal{T}$ = Binary_Decision_Tree($\mathcal{R}=\mathcal{C}$).
4: Evaluate the tree bottom-up. Start traversal from the root node $\mathcal{R}$.Call $\mathcal{M}$ = Evaluate($\mathcal{R}$)
5:return $\mathcal{M}$
6:function Binary_Decision_Tree($\mathcal{R}$)
7: Initialize node NN with value $\mathcal{R}$.
8: Count frequency of each element in $\mathcal{R}$.
9:$x=$ most frequent element.
10:if $x$ appears in more than one term in $\mathcal{R}$ then
11:  Call $\mathcal{R}_{with},\mathcal{R}_{without}=\textit{Split}(\mathcal{R},x)$
12:  Create a new node in the Binary Decision Tree on the left of node $N$, Call $N.left=\textit{Binary\_Decision\_Tree}(\mathcal{R}_{without})$
13:  Create a new node in the Binary Decision Tree on the right of node $N$, Call $N.right=\textit{Binary\_Decision\_Tree}(\mathcal{R}_{with})$
14:  Store $x$ in NN
15:else
16:  Node NN is a leaf node.
17:end if
18:function Split($\mathcal{R}$, x)
[…]
21:for each term $\tau\in\mathcal{R}$ do
22:  if $x\in\tau$ then
23:   $\mathcal{R}_{with}=\mathcal{R}_{with}\cup\{\tau\setminus\{x\}\}$
24:  else
25:   $\mathcal{R}_{without}=\mathcal{R}_{without}\cup\{\tau\}$
26:  end if
27:end for
28:return $\mathcal{C}_{with},\mathcal{C}_{without}$
29:function Evaluate(NN)
30: Start evaluation at the root node NN.
31:if NN is a leaf then
32:  return EvaluateSelf(N)(N)
33:else
34:  $\mathcal{L}eft=\textit{Evaluate}(N.left)$ if $N.left$ exists
35:  $\mathcal{R}ight=\textit{Evaluate}(N.right)$ if $N.right$ exists
36:  Combine $\mathcal{L}eft$ and $\mathcal{R}ight$ along with the most frequent element $x$ stored in node NN (from line 9) to get the updated value of current node NN.
37:  $N.\textit{evaluation}=\textit{Combine}(\mathcal{L}eft,\mathcal{R}ight,N.x)$
38:end if
39:return $N.\textit{evaluation}$
[…]
42:if $x$ exists then
43:  Add $\{x\}$ to $\mathcal{Q}$
44:end if
45:if $\mathcal{L}eft$ is empty then
46:  Add all sets in $\mathcal{R}ight$ to $\mathcal{Q}$
47:else if $\mathcal{R}ight$ is empty then
48:  Add all sets in $\mathcal{L}eft$ to $\mathcal{Q}$
49:else
50:  for each $l\in\mathcal{L}eft$ do
51:   for each $r\in\mathcal{R}ight$ do
52:    Add $l\cup r$ as a new element to $\mathcal{Q}$.
53:   end for
54:  end for
55:end if
56:return $\mathcal{Q}$
57:function EvaluateSelf(N)(N)
58: Leaf node with value $\mathcal{X}=\{X_{1},X_{2},\dots,X_{k}\}$
[…]
60:for $i=2$ to $k$ do
61:  $\mathcal{L}eafnode=\textit{Multiply}(\mathcal{L}eafnode,\mathcal{X}_{i})$
62:end for
63:if node has stored split element $x$ then
64:  $\mathcal{L}eafnode=\mathcal{L}eafnode\cup\{\{x\}\}$
65:end if
66:function Multiply(A,B)(A,B)
[…]
69:for each set $a\in A$ do
70:  if $a\cap B=\emptyset$ then
71:   for each $b\in B$ do
72:    Add $a\cup\{b\}$ to $\mathcal{Z}$
73:   end for
74:  else
75:   Add $a$ to $\mathcal{Z}$ {Absorb }
76:  end if
77:end for
The methodology used in Fast-MCS is shown in Alg. 2. It is also explained using the example topology in Fig. 3. Consider the source-destination pair STS-T. The MPS for this source-destination pair are first evaluated based on the method discussed in Sec. III-B. For the example in Fig. 3, STS-T pair has four MPS: SABTS-A-B-T, SADEFTS-A-D-E-F-T, SCDEFTS-C-D-E-F-T, and SCDBTS-C-D-B-T. After removing the source and destination nodes from the MPS, the MPS are represented as sets: $\{A,B\}$, $\{A,D,E,F\}$, $\{C,D,E,F\}$, and $\{C,D,B\}$. This is the first step shown in the illustrated example in Fig. 4, and detailed in Line 2 of Alg. 2.
Based on these sets, the most repeated element is selected, and the sets are split into two branches of a tree, (i) without the most repeated element on the left and (ii) with the most repeated element on the right, as shown in the function Binary_Decision_Tree in Line 6 of Alg. 2. The sets on the right are assumed to remove the most frequently occurring element as the common term. The branching on either side of the tree continues till there are no more repeated elements in the sets. In the aforementioned example, $D$ is the most repeated element. The only set without $D$, $\{A,B\}$, is written on the left branch. The other sets with $D$ are written on the right branch, assuming that $D$ is taken out as the common term, as $\{A,E,F\}$, $\{C,E,F\}$, and $\{C,B\}$ in the second step in Fig. 4. This procedure of splitting corresponds to the Split function in Line 18 of Alg. 2.
The set on the left $\{A.B\}$ cannot be further split. However, the sets on the right can be further split with $C$ as the most repeated element, giving the third step in Fig. 4 based on the Split function in Line 18 of Alg. 2. No further branching is possible, and hence the obtained branches are leaf nodes. After the recursive splitting operations, the branches must be simplified before combining. For example, the leaf node $\{E,F\}$, and $\{B\}$ must be multiplied based on the EvaluateSelf and Multiply functions in Lines 57 and 66 of Alg. 2, respectively. The two sets are multiplied and checked for absorption to remove any subsets. In this case, they are evaluated to be $\{BE,BF\}$. After the leaf nodes are evaluated, they must be combined using the Combine function in Line 40 of Alg. 2. For example, the branch with $\{A,E,F\}$ is combined with the aforementioned evaluation of the leaf node. The combination operation computes the union of each pair of elements from the two sets and checks for absorption, removing any resulting subsets if present. The combination operation is recursively repeated to identify the MCS for the given source-destination pair. For the example topology in Fig. 3 and the source-destination pair STS-T, the MCS are $\{(A,D),(A,C),(B,D),(B,E),(B,F)\}$.
III-E Differentiating Fast-MCS from Boole-Shannon expansion [26]
Though our Fast-MCS and the Boole-Shannon expansion used in [26] employ a recursive decomposition strategy, they differ fundamentally. The Boole-Shannon expansion is applied directly on the Boolean representation of the system’s success function (MPS in this work), manipulating logical expressions to derive MCS as the prime implicants of the complemented function. On the other hand, our Fast-MCS operates strictly on a set-theoretic representation of MPS and never constructs an explicit Boolean formula. Fast-MCS’s decision tree recursively partitions the sets into subsets that either contain or do not contain an element.
For large topologies, handling very long Boolean expressions, expansions, and absorptions is memory and time-consuming. However, Fast-MCS’s set operations are faster and more memory-efficient. This property is also clear from the results discussed in Sec. IV.
III-F Workflow
Fig. 5 shows the workflow adopted in this work. Given an input topology, first, MPS are evaluated for each source-destination pair. Then, the approaches to find MCS- combinatorial method, Boole-Shannon expansion, and our Fast-MCS are evaluated. The times required to find MPS and MCS using the different approaches are measured and then compared.
Figure 5: Workflow adopted in this work.
TABLE I: Topologies and their sources, evaluated in this study. $|V|$ and $|E|$ denote the number of nodes and edges in topology $G$.
$G$ $|V|$ $|E|$ $G$ $|V|$ $|E|$ $G$ $|V|$ $|E|$ $G$ $|V|$ $|E|$
Abilene [28] 11 14 Spain [29] 17 28 USA_26 [29] 26 42 India35 [29] 35 80
Polska [2] 12 18 Austria_24 [29] 24 50 Norway [29] 27 51 jonas-us-ca [29] 39 61
HiberniaUK [28] 13 13 Sweden [29] 25 29 Nobel_EU [2] 28 41 pioro40 [29] 40 89
Germany_17 [2] 17 26
IV Results
We evaluated Fast-MCS on an AMD Ryzen 5 3600 hexacore processor with 24 GB RAM running WSL2 [30] on 13 topologies shown in Table I. For each topology, each possible source-destination pair is considered. The workflow for Fast-MCS corresponds to that discussed in Fig. 5.
IV-A Fast-MCS vs. Combinatorial approach
[…]
Figure 6: Comparing the runtimes for Fast-MCS and the combinatorial approach [8, 17] (Sec. III-A). The blue and red bars correspond to finding MPS and MCS from MPS, respectively. Combined, the two bars together represent the total runtime for Fast-MCS. The orange bar corresponds to the combinatorial method. Fast-MCS outperforms the combinatorial method in every topology.
Figure 7: Comparing the runtimes for Fast-MCS and the Boole-Shannon expansion approach [26] (Sec. III-C) in the blue and orange bars, respectively. Boole-Shannon expansion approach offers slightly better performance than Fast-MCS for small topologies.
Figure 8: Comparing the runtimes for Fast-MCS and the Boole-Shannon expansion approach [26] (Sec. III-C) in the blue and orange bars, respectively. Fast-MCS provides significantly faster computation than Boole-Shannon expansion approach.
IV-B Fast-MCS vs.Boole-Shannon expansion
Fig. 7 shows Fast-MCS’s performance compared against the Shannon expansion method proposed in [26]. The Y axis shows the simulation time in seconds on a logarithmic scale for the corresponding topology in the X axis. The blue and orange bars correspond to the runtimes of Fast-MCS and Shannon expansion techniques, respectively. Shannon expansion [26] outperforms Fast-MCS in smaller topologies by less than 100 milliseconds. However, for larger topologies, Fast-MCS performs better by a few seconds. To further investigate this behavior, we tested both techniques on larger topologies as seen in Fig. 8. Here, Fast-MCS outperforms Shannon expansion by at least 80%. The runtime saved by using Fast-MCS in larger topologies far outweighs the runtime lost by using Fast-MCS in smaller topologies. Therefore, Fig. 8 shows that Fast-MCS is scalable and efficient in large topologies. Figs. 67 and 8 are plotted separately because the Y axis varies significantly for each. Plotting them together loses critical information portrayed in Fig. 8.
V Conclusions and Outlook
MCS (MCS) are crucial tools in evaluating a network’s resilience. In this work, we present our scalable, open-source tool, Fast-MCS, for efficiently obtaining MCS. We compared Fast-MCS’s performance against other methods, such as the combinatorial method and Shannon expansion, and show that Fast-MCS is faster than the other methods. Though the examples considered in this manuscript are communication network topologies, Fast-MCS is applicable to any system that can be represented as a graph. Currently, components are considered binary-valued, i.e., either working or unavailable. In the future, we intend to evaluate MCS as a time-varying function of components’ availabilities.